OpenStudy (anonymous):

f(x)= 3/x+6 g(x)=x^2-3 find[gxf](x)

OpenStudy (tkhunny):

Did you mean \(f(x) = \dfrac{3}{x} + 6\)? That's what you wrote.

OpenStudy (anonymous):

\[f(x)=\frac{ 3 }{x+6} \]

OpenStudy (tkhunny):

Parentheses are wonderful things: f(x) = 3/(x+6). You must see the difference. We seek g(f(x))?

OpenStudy (anonymous):

sorry... and yes that's what I'm looking for.

OpenStudy (tkhunny):

It's never a matter of "sorry". We're all learning. Can you evaluate g(7) = ?

OpenStudy (anonymous):

the questions wants me to just find the g(f(x)) not an input value

OpenStudy (tkhunny):

You didn't answer my question. Evaluate g(7).

OpenStudy (anonymous):

g(7)=46

OpenStudy (tkhunny):

How did you get that? Like this: g(7) = 7^2 - 3 = 49 - 3 = 46? You had g(x). When you moved to g(7), what did you do? Whatever you did, why would it be any different for g(4) or g(12) or g(Whale) or g(f(x))? Do EXACTLY the same thing with f(x) as you did with 7. EXACTLY!